OpenStudy (anonymous):

Prove that the roots of (a-b)^2x^2+2(a+b-2c)x+1=0 , a

OpenStudy (anonymous):

for the quadratic equation ax^2+bx+c=0 to have imaginary roots, the condition which is to be satisfied is b^2-4ac<0. use this condition to your equation and then find the condition for c.

OpenStudy (anonymous):

\[(a-b)^2x^2+2(a+b-2c)x+1=0\] \[b^2-4ac \implies \] \[(2(a+b-2c))^2-4(a-b)^2*1\]

OpenStudy (anonymous):

\[4(a^2+b^2+4c^2-2(ab-2ac-2bc)-a^2+2ab-b^2)\]

OpenStudy (anonymous):

\[4(4c^2-2ab+4ac+4bc+2ab)\]

OpenStudy (anonymous):

\[4(4c^2+4ac+4bc)\] \[16(c^2+c(a+b))\]

OpenStudy (anonymous):

this is real when \[b^2-4ac>0\] real roots \[c^2+c(a+b) \ge0\] \[c+a+b \ge 0\] \[c\ge a+b\]

OpenStudy (anonymous):

i mean \[a+b \ge c\]

OpenStudy (anonymous):

for imaginery \[a+b \le c\]

OpenStudy (anonymous):